OpenStudy (anonymous):

Write the equation of line that is perpendicular to y = 1/4x and passes through the point (2, -5)

OpenStudy (mathstudent55):

The slopes of perpendicular lines are negative reciprocals. That means the product of the slopes is -1. The line you are given, \(y = \dfrac{1}{4}x \) is in slope-intercept form, \(y = mx + b\), with slope m and y-intercept b. Its slope is \( \dfrac{1}{4} \). The slope of the perpendicular is \( - \dfrac{4}{1} = -4 \) \( y = -4x + b\) Now plug in the x and y coordinates of the point you are given into x and y in the equation above, and solve for b. That will give you the y-intercept.

OpenStudy (anonymous):

y=-4x-13?